OpenStudy (aimmebgood):

3a^(2) - 21a + 30 factor completely. The answer is 3(a - 2)(a-5) but thats not what I keep getting. how do you get to the answer?

OpenStudy (anonymous):

\[3 a^2-21 a+30=3\left(a^2-7 a+10\right)=3 (a-5) (a-2) \]

OpenStudy (aimmebgood):

where does the -5 and the -2 come from in the equation?

OpenStudy (anonymous):

quadratic formulas

OpenStudy (anonymous):

http://www.purplemath.com/modules/factquad.htm